OpenStudy (anonymous):

Complete the square by writing x^2+18x-19 in the form (x-h)^2+K

OpenStudy (anonymous):

the polynomial\[f(x) = x^2 + bx + c = (x + \frac{ b }{ 2 })^2 - \left( \frac{ b }{ 2 } \right)^2 + c\]